# KMM: Key Frame Mask Mamba for Extended Motion Generation

### v1

@@ Introduction
$\text{sim}_{ij}=\mathbf{T}_{i}^{\top}\mathbf{M}_{j}$ (12)
Then, the similarity is scaled by the temperature parameter $\tau$:
$\text{sim}_{ij}=\frac{\mathbf{T}_{i}^{\top}\mathbf{M}_{j}}{\tau}$ (13)
@@
$\mathcal{L}_{\text{contrast}}=\frac{1}{2}\left(\text{CrossEntropy}(\text{sim},\mathbf{y})+\text{CrossEntropy}(\text{sim}^{\top},\mathbf{y})\right)$
@@
TEACH $0.460^{\pm 0.000}$ $1.12^{\pm 0.00}$ $8.28^{\pm 0.00}$ $7.14^{\pm 0.00}$
TEACH w/o Spherical Linear Interpolation $\textbf{0.703}^{\pm 0.002}$ $1.71^{\pm 0.03}$ $8.18^{\pm 0.14}$ $\underline{3.43}^{\pm 0.01}$
TEACH $0.655^{\pm 0.002}$ $1.82^{\pm 0.02}$ $7.96^{\pm 0.11}$ $3.72^{\pm 0.01}$
PriorMDM $0.430^{\pm 0.000}$ $1.04^{\pm 0.00}$ $8.14^{\pm 0.00}$ $7.39^{\pm 0.00}$
PriorMDM w/ Trans. Emb $0.480^{\pm 0.000}$ $0.79^{\pm 0.00}$ $8.16^{\pm 0.00}$ $6.97^{\pm 0.00}$
PriorMDM w/ Trans. Emb & geo losses $0.450^{\pm 0.000}$ $0.91^{\pm 0.00}$ $8.16^{\pm 0.00}$ $7.09^{\pm 0.00}$
PriorMDM $0.596^{\pm 0.005}$ $3.16^{\pm 0.06}$ $7.53^{\pm 0.11}$ $4.17^{\pm 0.02}$
PriorMDM w/ PCCAT and APE $0.668^{\pm 0.005}$ $1.33^{\pm 0.04}$ $7.98^{\pm 0.12}$ $3.67^{\pm 0.03}$
MultiDiffusion $\underline{0.702}^{\pm 0.005}$ $1.74^{\pm 0.04}$ $\underline{8.37}^{\pm 0.13}$ $\underline{3.43}^{\pm 0.02}$
DiffCollage $0.671^{\pm 0.003}$ $1.45^{\pm 0.05}$ $7.93^{\pm 0.09}$ $3.71^{\pm 0.01}$
T2LM $0.589^{\pm 0.000}$ $0.66^{\pm 0.00}$ $8.99^{\pm 0.00}$ $3.81^{\pm 0.00}$
FlowMDM $\underline{0.702}^{\pm 0.004}$ $0.99^{\pm 0.04}$ $8.36^{\pm 0.13}$ $3.45^{\pm 0.02}$
Motion Mamba $0.490^{\pm 0.000}$ $0.76^{\pm 0.00}$ $\textbf{8.39}^{\pm 0.00}$ $4.97^{\pm 0.00}$
InfiniMotion $0.510^{\pm 0.000}$ $\underline{0.58}^{\pm 0.00}$ $8.67^{\pm 0.00}$ $4.89^{\pm 0.00}$
KMM (Ours) $0.666^{\pm 0.001}$ $\textbf{0.34}^{\pm 0.01}$ $8.67^{\pm 0.14}$ $\textbf{3.11}^{\pm 0.01}$
@@
Table 1: This table presents a comparison between our method and previous long motion generation techniques on the BABEL dataset (Punnakkal et al. 2021). The results show that our method outperforms the others, demonstrating superior performance. The right arrow $\rightarrow$ indicates that closer values to real motion are better. Bold and underline highlight the best and second-best results, respectively. Additionally, * denotes results reproduced by FlowMDM. For results with ±$\pm{0.000}$ or ±$\pm{0.00}$, the corresponding paper does not provide error bars.
Experiments
Datasets and Evaluation Matrices
@@
BABEL Dataset BABEL dataset (Punnakkal et al. 2021) is the go-to benchmark for long motion generation and has been widely adopted in previous extended motion generation work. Derived from AMASS (Mahmood et al. 2019), BABEL provides detailed annotations for extended motion sequences, with each segment linked to a specific textual annotation. The dataset includes a total of 10,881 motion sequences, consisting of 65,926 segments, each with its corresponding textual label.
@@
BABEL-D Dataset To evaluate the performance of text-motion alignment in extended motion generation methods, we introduce a new benchmark, BABEL-D. This benchmark is a subset of the BABEL test set and includes directional conditions with keywords such as left and right. This also represents the more challenging subset of BABEL. The BABEL-D dataset contains a total of 560 motion segments, enabling us to demonstrate improved alignment between generated motion and given text queries. We then evaluate our method’s performance on BABEL-D and compare it with other state-of-the-art extended motion generation approaches.
Why not HumanML3D? Despite HumanML3D (Guo et al. 2022) is an important dataset for text-to-motion generation, however it’s not suitable for extended motion generation since it only contains motion segments which is less than 196 frames. This limitation restrict the potential for evaluating long motion generation and thereby cannot reflect the real long sequence generation capability of these methods. However, the quality of text annotations in HumanML3D still benefits text-to-motion generation in general, including long motion generation. To showcase our method’s generalizability, our qualitative visualization combines different text queries from HumanML3D for long motion generation.
@@
Evaluation Matrices For our experiments, we adopted the quantitative evaluation matrices for text-to-motion generation originally introduced by T2M (Guo et al. 2022) and later used in long motion generation studies (Shafir et al. 2023; Barquero, Escalera, and Palmero 2024; Zhang et al. 2024b). These include: (1) Frechet Inception Distance (FID), which measures overall motion quality by assessing the distributional difference between the high-level features of generated and real motions; (2) R-precision; (3) MultiModal Distance, both of which evaluate the semantic alignment between the input text and generated motions; and (4) Diversity, which calculates the variance in features extracted from the motions. The main difference in our evaluation compared to previous long motion generation methods is that we treat complex text queries as a whole, rather than generating separate motion segments then interpolating transitions between them. This eliminates the need to evaluate transition interpolation and represents a significant advancement in long motion generation.
Comparative Studies
@@
Evaluation on BABEL To evaluate the performance of our KMM on extended motion generation, we trained and evaluated it on the BABEL dataset. The results, as shown in Tables 1, indicate that our method significantly outperforms previous text-to-motion generation approaches specifically designed for long-sequence motion generation. All experiments were conducted with a batch size of 256 for RVQ-VAE, which utilized 6 quantization layers, and a batch size of 64 for mask bidirectional Mamba. These experiments were carried out on a single Intel Xeon Platinum 8360Y CPU at 2.40GHz, paired with a single NVIDIA A100 40G GPU and 32GB of RAM.
@@
Evaluation on BABEL-D To quantitatively demonstrate the advantages of our proposed text-motion alignment method in addressing directional instructions, we conducted comprehensive experiments on the newly introduced BABEL-D benchmark. The results have shown in Table 2. Compared to previous state-of-the-art methods, our approach significantly outperforms other extended motion generation techniques, indicating a stronger alignment between text and motion.
@@
PriorMDM $0.334^{\pm 0.015}$ $6.82^{\pm 0.76}$ $7.27^{\pm 0.33}$ $7.44^{\pm 0.12}$
KMM w/o Alignment $\underline{0.484}^{\pm 0.007}$ $\underline{5.50}^{\pm 0.15}$ $\textbf{8.44}^{\pm 0.15}$ $\underline{3.48}^{\pm 0.03}$
KMM (Ours) $\textbf{0.538}^{\pm 0.009}$ $\textbf{3.86}^{\pm 0.14}$ $\underline{8.04}^{\pm 0.14}$ $\textbf{2.72}^{\pm 0.03}$
Table 2: This table compares our method with previous long motion generation techniques on the BABEL-D benchmark. The results demonstrate that our method excels in handling directional instructions, highlighting the advantages of our proposed text-motion alignment approach. The right arrow $\rightarrow$ indicates that closer values to real motion are better. Bold and underline highlight the best and second-best results, respectively.
Ablation Study
To further evaluate different aspects of our method’s impact on overall performance, we conducted various ablation studies on the BABEL dataset, as shown in Table 3. The results show that our approach substantially outperforms other masking strategies, including random masking, KMeans (Lloyd 1982), and GMM (Reynolds et al. 2009) key frame selection. Additionally, our proposed text-motion alignment framework greatly improves the model’s ability to understand complex text queries, leading to better-aligned motion sequences.
@@
KMM w/o Alignment $\textbf{0.671}^{\pm 0.001}$ $\underline{0.40}^{\pm 0.01}$ $\textbf{8.57}^{\pm 0.05}$ $\underline{3.21}^{\pm 0.01}$
KMM (Ours) $\underline{0.666}^{\pm 0.001}$ $\textbf{0.34}^{\pm 0.01}$ $\underline{8.67}^{\pm 0.14}$ $\textbf{3.11}^{\pm 0.01}$
Table 3: This table illustrates the ablation results from different aspects of the proposed method. The results show that both the key frame masking strategy and text-motion alignment contribute to the overall performance. The right arrow $\rightarrow$ indicates that closer values to real motion are better. Bold and underline highlight the best and second-best results, respectively.
Conclusion
In conclusion, our study addresses two significant challenges in extended motion generation: memory decay in long sequence generation and weak text-motion alignment. Our proposed method, KMM, presents innovative solutions that significantly advance the field. Our density-based key frame selection and masking strategy enhances Mamba’s ability to focus on critical actions within long motion sequences, effectively mitigating the memory decay problem. Additionally, our robust contrastive learning paradigm improves text-motion alignment, enabling more accurate motion generation for complex and directional text queries. Furthermore, the development of the BABEL-D benchmark provides a valuable resource for evaluating text-motion alignment in extended motion generation, specifically focused on directional instructions. This new dataset, alongside our comprehensive experiments on the BABEL dataset, underscores our commitment to advancing the field of motion generation across various domains.
KMM Supplementary Materials

### v2

@@ 3.3. Text-Motion Alignment
(12) $\text{sim}_{ij}=\mathbf{T}_{i}^{\top}\mathbf{M}_{j}.$
Then, the similarity is scaled by a learnable temperature parameter $\tau$:
(13) $\text{sim}_{ij}=\frac{\mathbf{T}_{i}^{\top}\mathbf{M}_{j}}{\tau}.$
@@
where the coefficient $\lambda$ is set to 0.5.
@@
Table 1. Comparison on BABEL (Punnakkal et al., 2021). The right arrow $\rightarrow$ indicates that closer values to real motion are better. Bold and underline highlight the best and second-best results, respectively. Additionally, * denotes results reproduced by FlowMDM. SLI denotes spherical linear interpolation. For results with ±$\pm{0.000}$ or ±$\pm{0.00}$, the corresponding paper does not provide error bars.
Subsequence Transition
[…]
Ground Truth $0.715^{\pm 0.003}$ $0.00^{\pm 0.00}$ $8.42^{\pm 0.15}$ $3.36^{\pm 0.00}$ $0.00^{\pm 0.00}$ $6.20^{\pm 0.06}$ $0.02^{\pm 0.00}$ $0.00^{\pm 0.00}$
TEACH (Athanasiou et al., 2022) $0.460^{\pm 0.000}$ $1.12^{\pm 0.00}$ $8.28^{\pm 0.00}$ $7.14^{\pm 0.00}$ $7.93^{\pm 0.00}$ $6.53^{\pm 0.00}$
TEACH w/o SLI (Athanasiou et al., 2022) $\mathbf{0.703}^{\pm 0.002}$ $1.71^{\pm 0.03}$ $8.18^{\pm 0.14}$ $\underline{3.43}^{\pm 0.01}$ $3.01^{\pm 0.04}$ $\mathbf{6.23}^{\pm 0.05}$ $1.09^{\pm 0.00}$ $2.35^{\pm 0.01}$
TEACH (Athanasiou et al., 2022) $0.655^{\pm 0.002}$ $1.82^{\pm 0.02}$ $7.96^{\pm 0.11}$ $3.72^{\pm 0.01}$ $3.27^{\pm 0.04}$ $6.14^{\pm 0.06}$ ${0.07}^{\pm 0.00}$ ${0.44}^{\pm 0.00}$
PriorMDM (Shafir et al., 2023) $0.430^{\pm 0.000}$ $1.04^{\pm 0.00}$ $8.14^{\pm 0.00}$ $7.39^{\pm 0.00}$ $3.45^{\pm 0.00}$ $7.19^{\pm 0.00}$
PriorMDM w/ Trans. Emb (Shafir et al., 2023) $0.480^{\pm 0.000}$ $0.79^{\pm 0.00}$ $8.16^{\pm 0.00}$ $6.97^{\pm 0.00}$ $7.23^{\pm 0.00}$ $6.41^{\pm 0.00}$
PriorMDM w/ Trans. Emb & geo losses (Shafir et al., 2023) $0.450^{\pm 0.000}$ $0.91^{\pm 0.00}$ $8.16^{\pm 0.00}$ $7.09^{\pm 0.00}$ $6.05^{\pm 0.00}$ $6.57^{\pm 0.00}$
PriorMDM (Shafir et al., 2023) $0.596^{\pm 0.005}$ $3.16^{\pm 0.06}$ $7.53^{\pm 0.11}$ $4.17^{\pm 0.02}$ $3.33^{\pm 0.06}$ $\underline{6.16}^{\pm 0.05}$ $0.28^{\pm 0.00}$ $1.04^{\pm 0.01}$
PriorMDM w/ PCCAT and APE (Shafir et al., 2023) $0.668^{\pm 0.005}$ $1.33^{\pm 0.04}$ $7.98^{\pm 0.12}$ $3.67^{\pm 0.03}$ $3.15^{\pm 0.05}$ $6.14^{\pm 0.07}$ $0.17^{\pm 0.00}$ $0.64^{\pm 0.01}$
MultiDiffusion (Bar-Tal et al., 2023) $\underline{0.702}^{\pm 0.005}$ $1.74^{\pm 0.04}$ $\underline{8.37}^{\pm 0.13}$ $\underline{3.43}^{\pm 0.02}$ $6.56^{\pm 0.12}$ $5.72^{\pm 0.07}$ $0.18^{\pm 0.00}$ $0.68^{\pm 0.00}$
DiffCollage (Zhang et al., 2023b) $0.671^{\pm 0.003}$ $1.45^{\pm 0.05}$ $7.93^{\pm 0.09}$ $3.71^{\pm 0.01}$ $4.36^{\pm 0.09}$ $6.09^{\pm 0.08}$ $0.19^{\pm 0.00}$ $0.84^{\pm 0.01}$
T2LM (Lee et al., 2024) $0.589^{\pm 0.000}$ $0.66^{\pm 0.00}$ $8.99^{\pm 0.00}$ $3.81^{\pm 0.00}$
FlowMDM (Barquero et al., 2024) $\underline{0.702}^{\pm 0.004}$ $0.99^{\pm 0.04}$ $8.36^{\pm 0.13}$ $3.45^{\pm 0.02}$ $\underline{2.61}^{\pm 0.06}$ $6.47^{\pm 0.05}$ $\mathbf{0.06}^{\pm 0.00}$ $\underline{0.13}^{\pm 0.00}$
Motion Mamba (Zhang et al., 2024c) $0.490^{\pm 0.000}$ $0.76^{\pm 0.00}$ $\mathbf{8.39}^{\pm 0.00}$ $4.97^{\pm 0.00}$
InfiniMotion (Zhang et al., 2024b) $0.510^{\pm 0.000}$ $\underline{0.58}^{\pm 0.00}$ $8.67^{\pm 0.00}$ $4.89^{\pm 0.00}$
KMM (Ours) $0.666^{\pm 0.001}$ $\mathbf{0.34}^{\pm 0.01}$ $8.67^{\pm 0.14}$ $\mathbf{3.11}^{\pm 0.01}$ $\mathbf{1.37}^{\pm 0.04}$ ${5.96}^{\pm 0.09}$ $\underline{0.08}^{\pm 0.00}$ $\mathbf{0.10}^{\pm 0.00}$
4. Experiments
4.1. Datasets and Evaluation Matrices
BABEL Dataset.
@@ 4.1. Datasets and Evaluation Matrices
BABEL (Punnakkal et al., 2021) is the go-to benchmark for long motion generation and has been widely adopted in previous extended motion generation work. Derived from AMASS (Mahmood et al., 2019), BABEL provides both frame-level and motion annotations for extended motion sequences. The dataset includes a total of 10,881 motion sequences, consisting of 65,926 segments, each with its corresponding textual label.
BABEL-D Dataset.
@@
To evaluate the performance of text-motion alignment in extended motion generation methods, we introduce a new benchmark, BABEL-D. This benchmark is a subset of the BABEL test set and includes directional conditions with keywords such as left and right. This also represents the more challenging subset of BABEL. The BABEL-D dataset contains a total of 560 motion segments, enabling us to demonstrate improved alignment between generated motion and given text queries. We then evaluate our method’s performance on BABEL-D and compare it with other state-of-the-art extended motion generation approaches.
HumanML3D Dataset.
HumanML3D (Guo et al., 2022) is the go-to dataset for text-to-motion generation, including 14,616 motions with text descriptions. Despite the maximum length of HumanML3D being only 196 frames, we also evaluate our method on this dataset to demonstrate its generalizability.
Evaluation Matrices.
@@
For our experiments, we adopted the quantitative evaluation matrices for text-to-motion generation originally introduced by T2M (Guo et al., 2022) and later used in long motion generation studies (Shafir et al., 2023; Barquero et al., 2024; Zhang et al., 2024b). These include: (1) Frechet Inception Distance (FID), which measures overall motion quality by assessing the distributional difference between the high-level features of generated and real motions; (2) R-precision; (3) MultiModal Distance, both of which evaluate the semantic alignment between the input text and generated motions; and (4) Diversity, which calculates the variance in features extracted from the motions. For transition evaluation, we adopt two metrics from FlowMDM (Barquero et al., 2024). Peak Jerk (PJ) captures the maximum jerk across joints to identify abrupt changes. However, it may favor overly smoothed transitions. To address this, we also include Area Under the Jerk (AUJ), which quantifies deviations from average jerk using L1-norm differences.
4.2. Comparative Study
Evaluation on BABEL.
@@ 4.2. Comparative Study
To evaluate the performance of our KMM on extended motion generation, we trained and evaluated it on the BABEL dataset. The results, as shown in Tables 1, indicate that our method significantly outperforms previous text-to-motion generation approaches specifically designed for long-sequence motion generation. All experiments were conducted with a batch size of 256 for VQ-VAE, which utilized 6 quantization layers, and a batch size of 64 for mask bidirectional Mamba. These experiments were carried out on a single Intel Xeon Platinum 8360Y CPU at 2.40GHz, paired with a single NVIDIA A100 40G GPU and 32GB of RAM.
Evaluation on BABEL-D.
@@
To quantitatively demonstrate the advantages of our proposed text-motion alignment method in addressing directional instructions, we conducted comprehensive experiments on the newly introduced BABEL-D benchmark. The results have shown in Table 2. Compared to previous state-of-the-art methods, our approach significantly outperforms other extended motion generation techniques, indicating a stronger alignment between text and motion.
Table 2. Comparison on BABEL-D. The right arrow $\rightarrow$ indicates that closer values to real motion are better. Bold and underline highlight the best and second-best results, respectively.
@@
PriorMDM (Shafir et al., 2023) $0.334^{\pm 0.015}$ $6.82^{\pm 0.76}$ $7.27^{\pm 0.33}$ $7.44^{\pm 0.12}$
FlowMDM (Barquero et al., 2024) $\underline{0.535}^{\pm 0.010}$ $\underline{1.45}^{\pm 0.07}$ $\underline{8.09}^{\pm 0.09}$ $\underline{2.87}^{\pm 0.03}$
KMM w/o Alignment ${0.484}^{\pm 0.007}$ ${5.50}^{\pm 0.15}$ $\mathbf{8.44}^{\pm 0.15}$ ${3.48}^{\pm 0.03}$
KMM (Ours) $\mathbf{0.538}^{\pm 0.009}$ $\mathbf{0.62}^{\pm 0.03}$ ${8.04}^{\pm 0.14}$ $\mathbf{2.72}^{\pm 0.03}$
Evaluation on HumanML3D.
We conducted experiments on HumanML3D (Guo et al., 2022) and compared our results with previous state-of-the-art long-motion methods. The results are presented in the Table 3, indicating that our KMM method significantly outperforms previous long-motion methods and demonstrates strong generalizability across multiple datasets.
Table 3. Comparison on HumanML3D (Guo et al., 2022). The right arrow $\rightarrow$ indicates that closer values to real motion are better. Bold and underline highlight the best and second-best results, respectively.
[…]
Ground Truth $0.796^{\pm 0.004}$ $0.00^{\pm 0.00}$ $9.34^{\pm 0.08}$ $2.97^{\pm 0.01}$
MultiDiffusion (Bar-Tal et al., 2023) $0.629^{\pm 0.002}$ $1.19^{\pm 0.03}$ $\underline{9.38}^{\pm 0.08}$ $4.02^{\pm 0.01}$
DiffCollage (Zhang et al., 2023b) $0.615^{\pm 0.005}$ $1.56^{\pm 0.04}$ $8.79^{\pm 0.08}$ $4.13^{\pm 0.02}$
PriorMDM (Shafir et al., 2023) $0.590^{\pm 0.000}$ $0.60^{\pm 0.00}$ $9.50^{\pm 0.00}$ $5.61^{\pm 0.00}$
FlowMDM (Barquero et al., 2024) $\underline{0.685}^{\pm 0.004}$ $\underline{0.29}^{\pm 0.01}$ ${9.58}^{\pm 0.12}$ $\underline{3.61}^{\pm 0.01}$
KMM (Ours) $\mathbf{0.787}^{\pm 0.005}$ $\mathbf{0.15}^{\pm 0.01}$ $\mathbf{9.37}^{\pm 0.01}$ $\mathbf{3.08}^{\pm 0.02}$
Figure 5. The figure presents some qualitative visualization results of KMM. The text prompts are sourced and combined from HumanML3D (Guo et al., 2022) and BABEL (Punnakkal et al., 2021). The number within the brackets indicates our ability to condition the generated motion on a specific length, dynamically producing motion of the desired duration. The visualizations showcase KMM’s superior performance in generating robust and diverse motions that align closely with lengthy and complex text queries.
4.3. Ablation Study
To further evaluate different aspects of our method’s impact on overall performance, we conducted various ablation studies on the BABEL (Punnakkal et al., 2021), as shown in Table 4. The results show that our approach substantially outperforms other masking strategies, including random masking, KMeans (Lloyd, 1982), and GMM (Reynolds et al., 2009) key frame selection. Additionally, our proposed text-motion alignment framework greatly improves the model’s ability to understand complex text queries, leading to better-aligned motion sequences.
Table 4. Masking strategies. The right arrow $\rightarrow$ indicates that closer values to real motion are better. Bold and underline highlight the best and second-best results, respectively.
@@ 4.3. Ablation Study
KMM w/ KMeans $\underline{0.661}^{\pm 0.001}$ $0.43^{\pm 0.07}$ $\mathbf{8.38}^{\pm 0.09}$ $3.16^{\pm 0.01}$
KMM w/ GMM ${0.659}^{\pm 0.002}$ $\underline{0.40}^{\pm 0.01}$ $\underline{8.30}^{\pm 0.26}$ $\underline{3.12}^{\pm 0.01}$
KMM w/o Alignment $\underline{0.661}^{\pm 0.001}$ $\underline{0.40}^{\pm 0.01}$ ${8.57}^{\pm 0.05}$ ${3.21}^{\pm 0.01}$
KMM (Ours) $\mathbf{0.666}^{\pm 0.001}$ $\mathbf{0.34}^{\pm 0.01}$ ${8.67}^{\pm 0.14}$ $\mathbf{3.11}^{\pm 0.01}$
We also conducted comprehensive ablation studies on the key frame masking ratio and the coefficient $\lambda$ in the contrastive loss for text-motion alignment on BABEL (Punnakkal et al., 2021). The results are shown in Tables 5 and 6, demonstrating that our method is robust across different hyperparameter settings.
5. Qualitative Evaluation
Table 5. Masking ratio. The right arrow $\rightarrow$ indicates that closer values to real motion are better. Bold highlights the best results.
[…]
15% $0.661^{\pm 0.001}$ $0.69^{\pm 0.01}$ ${8.33}^{\pm 0.15}$ $3.27^{\pm 0.01}$
$\mathbf{30\%}$ (Ours) $\mathbf{0.666}^{\pm 0.001}$ $\mathbf{0.34}^{\pm 0.01}$ $\mathbf{8.67}^{\pm 0.14}$ $\mathbf{3.11}^{\pm 0.01}$
50% $0.063^{\pm 0.003}$ $0.41^{\pm 0.01}$ $8.79^{\pm 0.01}$ $3.26^{\pm 0.01}$
Table 6. Coefficient $\lambda$. The right arrow $\rightarrow$ indicates that closer values to real motion are better. Bold highlights the best results.
[…]
0.3 $0.667^{\pm 0.003}$ $0.40^{\pm 0.01}$ $8.64^{\pm 0.08}$ $3.25^{\pm 0.01}$
$\mathbf{0.5}$ (Ours) $\mathbf{0.666}^{\pm 0.001}$ $\mathbf{0.34}^{\pm 0.01}$ $\mathbf{8.67}^{\pm 0.14}$ $\mathbf{3.11}^{\pm 0.01}$
0.7 $0.680^{\pm 0.003}$ $0.39^{\pm 0.01}$ $8.83^{\pm 0.04}$ $3.25^{\pm 0.01}$
To further evaluate our method qualitatively, we compared KMM with TEACH (Athanasiou et al., 2022), PriorMDM (Shafir et al., 2023), and FlowMDM (Barquero et al., 2024) by generating a diverse set of prompts, randomly extracted and combined from the HumanML3D (Guo et al., 2022) and BABEL (Punnakkal et al., 2021) test sets. Figure 4 shows three of these comparisons, demonstrating that our method significantly outperforms others in handling complex text queries and generating more accurate corresponding motions. Moreover, to further demonstrate the robustness and diversity of motions generated by our KMM, we produced 15 additional sequences using text prompts randomly extracted and combined from the HumanML3D (Guo et al., 2022) and BABEL (Punnakkal et al., 2021) test sets. The results, shown in figure 5, highlight superior performance in generating robust and diverse motions that closely align with lengthy and complex text queries.
6. Conclusion
In conclusion, our study addresses two significant challenges in extended motion generation: memory limitations of Mamba’s hidden state for long sequence generation and weak text-motion alignment. Our proposed method, KMM, presents innovative solutions that significantly advance the field. Our density-based key frame selection and masking strategy enhances Mamba’s ability to focus on critical actions within long motion sequences, effectively mitigating the memory limitation problem. Additionally, our robust contrastive learning paradigm improves text-motion alignment, enabling more accurate motion generation for complex and directional text queries. Furthermore, the development of the BABEL-D benchmark provides a valuable resource for evaluating text-motion alignment in extended motion generation, specifically focused on directional instructions. This new dataset, alongside our comprehensive experiments on the BABEL dataset, underscores our commitment to advancing the field of motion generation across various domains.
Appendix
Figure 6. The figure shows the user study interface where 50 participants evaluated motion sequences generated by TEACH, PriorMDM, FlowMDM, and KMM, focusing on text-motion alignment, robustness, diversity, and usability. The text prompt are randomly extracted and combined from the HumanML3D (Guo et al., 2022) and BABEL (Punnakkal et al., 2021) test set.